OpenStudy (anonymous):

the number of years a person born in the united states is expected to live can be approximated by the equation y = 0.189x + 70.8, where x is the number of years since 1970. solve this equation for x. use this new equation life expectancy will be 80.2 years.

OpenStudy (anonymous):

y = 0.189x + 70.8 y - 70.8 = 0.189x (y - 70.8)/0.189 = x When y = 80.2, x = (80.2 - 70.8)/0.182 = 49.7 The year, then, would be 1970 +49.7 = 2019.7